OpenStudy (anonymous):

find an equation of the line having the given line and containing the given point. m=8/9, (2,-7)

OpenStudy (amistre64):

which form of the equation you need?

OpenStudy (anonymous):

y= mx+b (-7) = (8/9) ( 2) + b Solve for b and plug back into equation

OpenStudy (anonymous):

opps, yes, the equation of the line is y=__________

OpenStudy (amistre64):

y+7 = \(\frac{8}{9}\)(x-2) .... is slope point form :)

OpenStudy (amistre64):

y = (8/9)x -(8(2)/9) -7 gets you there then

OpenStudy (amistre64):

y = (8/9)x -(8(2)/9) -7(9)/9 y = (8/9)x+(-16-63)/9 y = (8/9)x - 79/9

OpenStudy (anonymous):

thank you thank you!

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

i cant believe im 25 and cant figure this math out! absurd!

OpenStudy (amistre64):

thats ok, wait till your my age and cant remember it anymore ;)